# Crowns as retracts

### v5

@@ 1 Introduction
With the exception of Theorem 4, all our results are valid also for infinite posets without infinite chains. Even if the maximal sets addressed in Definition 3 do not exist, everything remains correct if we replace ${\cal F}(P)$ by ${\cal F}_{0}(P)$ in Definition 4 of the multigraph $\mathfrak{F}(P)$.
@@ 2.1 Notation and basic terms
Due to $[a,v]_{P}\cap[b,w]_{P}=[a,w]_{P}\cap[b,v]_{P}$, this definition is independent of the choice of the disjoint edges of $C$. As indicated by the addendum “in $P$”, a 4-crown is proper or improper only with respect to the poset it belongs to, not as an induced sub-poset as such. However, because the ambient poset is always clear in what follows, we omit explicit reference to it in most cases. In parts of the early literature [5, 7, 9, 11, 13], improper 4-crowns are not regarded as crowns.
@@
For a homomorphism $f$ between posets or graphs, we denote by $f|_{Z}$ the pre-restriction of $f$ to a subset $Z$ of its domain. An order homomorphism $r:P\rightarrow P$ is called a retraction of the poset $P$ iff $r$ is idempotent, and an induced sub-poset $R$ of $P$ is called a retract of $P$ iff a retraction $r:P\rightarrow P$ exists with $r[X]=R$. For the sake of simplicity, we identify $r$ with its post-restriction and write $r:P\rightarrow R$. A poset $P$ has the fixed point property iff every retract of $P$ has the fixed point property [15, Theorem 4.8]. An improper 4-crown cannot be a retract of $P$.
@@ 2.2 Results about homomorphisms
There exists a surjective homomorphism $f:P\rightarrow C$ iff there exists a partition $A,B$ of $L(P)$ and a partition $V,W$ of $U(P)$ fulfilling
$\begin{split}&\>\textrm{If }\>F\subseteq E(P)\>\textrm{is a four-element set with}\>F\cap N\not=\emptyset\>\textrm{ for all}\>\\ &\;N\in\{A,B,V,W\},\>\textrm{then}\>F\>\textrm{is not an improper 4-crown of }\>P.\end{split}$ (3)
Moreover, in the case of $C\subseteq E(P)$, $C$ is a retract of $P$ iff partitions $A,B$ of $L(P)$ and $V,W$ of $U(P)$ exist which fulfill (3) and in which the points $a,b,v,w$ all belong to different sets.
@@
Only direction $\Leftarrow$” has to be proven. Assume that there exists a partition $A,B$ of $L(P)$ and a partition $V,W$ of $U(P)$ fulfilling (3). Let $f:E(P)\rightarrow C$ be a strict surjective homomorphism induced by these partitions.
For $x\in X$, we define the sets $\alpha(x)$ and $\beta(x)$ as in Proposition 1. Then $\alpha(x)\subseteq\{a,b\}$ and $\beta(x)\subseteq\{v,w\}$. Assume $\alpha(x)=\{a,b\}$ and $\beta(x)=\{v,w\}$. Then $x\in M(P)$, and there exist points $a^{\prime}\in f^{-1}(a)\cap{\downarrow\,}_{P}x$, $b^{\prime}\in f^{-1}(b)\cap{\downarrow\,}_{P}x$, $v^{\prime}\in f^{-1}(v)\cap{\uparrow\,}_{P}x$, and $w^{\prime}\in f^{-1}(w)\cap{\uparrow\,}_{P}x$. But then $F:=\{a^{\prime},b^{\prime},v^{\prime},w^{\prime}\}\subseteq E(P)$ is an improper 4-crown of $P$ with $F\cap N\not=\emptyset$ for all $N\in\{A,B,V,W\}$ in contradiction to (3). At least one of the sets $\alpha(x)$ and $\beta(x)$ is thus a singleton, and Proposition 1 yields the existence of an extension $g:P\rightarrow C$ of $f$.
In the case of $C\subseteq E(P)$ with $a,b,v,w$ all belonging to different sets of the partitions, the mapping $\pi\circ g$ is a retraction for an appropriate automorphism $\pi$ of $C$.
@@ 3.1 The multigraphs $\mathfrak{C}$ and $\mathfrak{F}(P)$
We call a homomorphism $\phi:\mathfrak{F}(P)\rightarrow\mathfrak{C}$ separating iff there exist $x,x^{\prime}\in L(P)$ with $x\not=x^{\prime}$ and $y,y^{\prime}\in U(P)$ with $y\not=y^{\prime}$ fulfilling
for all $\>\textrm{for all }\>F\in{\cal F}(P)\>\textrm{ :}\>$$\begin{cases}\;x\in F&\quad\Rightarrow\quad\phi(F)\notin{\cal A},\\ \;x^{\prime}\in F&\quad\Rightarrow\quad\phi(F)\notin{\cal B},\\ \;y\in F&\quad\Rightarrow\quad\phi(F)\notin{\cal V},\\ \;y^{\prime}\in F&\quad\Rightarrow\quad\phi(F)\notin{\cal W},\end{cases}$ (4)
and we call it $\{x,x^{\prime},y,y^{\prime}\}$-separating, if the set $\{x,x^{\prime},y,y^{\prime}\}$ is a 4-crown in $E(P)$.
@@
Let $P$ be a poset with at least two minimal points and at least two maximal points.
  1. 1.
    The existence of a surjective homomorphism from $P$ onto $C$ is equivalent to the existence of a separating homomorphism from $\mathfrak{F}(P)$ to $\mathfrak{C}$.
  2. 2.
    In the case of $C\subseteq E(P)$, the 4-crown $C$ is a retract of $P$ iff a $C$-separating homomorphism from $\mathfrak{F}(P)$ to $\mathfrak{C}$ exists.
@@
1. $\Rightarrow$”: Assume that a surjective homomorphism from $P$ onto $C$ exists. According to Lemma 1, there exists a partition $A,B$ of $L(P)$ and a partition $V,W$ of $U(P)$ fulfilling (3). With $g:E(P)\rightarrow C$ being a strict surjective homomorphism induced by these partitions, it is easily seen that the mapping
@@
is a homomorphism from $\mathfrak{F}(P)$ to $\mathfrak{C}$. There exists an $a^{\prime}\in L(P)$ with $g(a^{\prime})=a$, thus $a^{\prime}\in F\Rightarrow\phi(F)\notin{\cal B}$. Proceeding with pre-images of the other elements of $C$, we see that $\phi$ is separating.
$\Leftarrow$”: Let $\phi:\mathfrak{F}(P)\rightarrow\mathfrak{C}$ be a separating homomorphism. We define
@@
According to (4), we have $A_{0}\not=L(P)\not=B_{0}$. Because the set $L(P)$ contains at least two elements, we can distribute the elements of $L(P)\setminus(A_{0}\cup B_{0})$ arbitrarily on the sets $A_{0}$ and $B_{0}$ in such a way that the resulting sets $A\supseteq A_{0}$ and $B\supseteq B_{0}$ are both non-empty. The sets $A,B$ form a partition of $L(P)$. In the same way, we construct a partition $V,W$ of $U(P)$ with $V_{0}\subseteq V$ and $W_{0}\subseteq W$.
Let $G\in{\cal F}(P)$. There exists a set ${\cal N}\in\{{\cal A},{\cal B},{\cal V},{\cal W}\}$ with $\phi(G)\in{\cal N}$, say, ${\cal N}={\cal A}$. Then $L(P)\cap G\subseteq A_{0}$, thus $G\cap B=\emptyset$. In the same way, we see that also for the other choices of ${\cal N}$ there exists a set $N\in\{A,B,V,W\}$ with $G\cap N=\emptyset$, and Lemma 1 delivers the existence of a surjective homomorphism from $P$ onto $C$.
2. $\Rightarrow$”: Proven as for part 1, using the addendum in Lemma 1.
$\Leftarrow$”: Let $\phi:\mathfrak{F}(P)\rightarrow\mathfrak{C}$ be a $C$-separating homomorphism. In particular, $C$ is a proper 4-crown in $E(P)$.
We proceed as in the proof of part 1. Because $\phi$ is $C$-separating, we cannot have $a,b\in A_{0}$ or $a,b\in B_{0}$. Regardless to which of the three disjoint sets $A_{0},B_{0}$, or $L(P)\setminus(A_{0}\cup B_{0})$ the points $a$ and $b$ belong, we can ensure that after building the partition $A,B$ of $L(P)$, one of them is contained in $A$ and the other one in $B$. In the same way we build a partition $V,W$ of $U(P)$ in which $v$ and $w$ belong to different sets.
As in the proof of part 1, we see that the partitions $A,B$ and $V,W$ fulfill (3) ($C$ is a proper 4-crown!). Now the addendum in Lemma 1 yields a retraction of $P$ to $C$.
@@
Let $Z:=\left(\cup\;{\cal F}(P)\right)\cup{\cal M}(P)\cup C$ and $Q:=P|_{Z}$. Because of $\mathfrak{F}(P)=\mathfrak{F}(Q)$ and Theorem 1, the 4-crown $C$ is a retract of $P$ iff it is a retract of $Q$. The elements of $X\setminus Z$ are thus irrelevant for $C$ being a retract of $P$ or not. And indeed, for $x\in M(P)\setminus Z$, one of the sets $L(P)\cap{\downarrow\,}x$ or $U(P)\cap{\uparrow\,}x$ is a singleton, and according to [15, Ex. 4-5], the point $x$ can be removed by a series of I-retractions. By doing so for all points in $M(P)\setminus Z$, we arrive at the poset $P^{\prime}:=P|_{E(P)\cup Z}$. The poset $P^{\prime}|_{Z}$ contains $C$, the improper 4-crowns of $P$ and their inner, and the poset $P^{\prime}|_{E(P)\setminus Z}$ has height one or is an antichain. Intuitively, it is plausible that a homomorphism mapping $P^{\prime}|_{Z}$ onto $C$ can be extended to the total poset $P^{\prime}$.
@@ 4 Additional results about crowns as retracts
For the next theorem, we need a slight variant of Rival [11, pp.  118-119], [15, Lemma 4.35]. The proof is a copy of the short proof of [15, Lemma 4.35]:
@@
We conclude that for $P$ having the fixed point property, every edge of every crown in $E(P)$ must belong to an improper 4-crown in $E(P)$. (Examples are the rightmost poset in Figure 1 and the posets in Figure 3.) For a poset $P$ of height two, this necessary condition can be sharpened: every edge of every crown in $E(P)$ must belong to an hourglass-crown, i.e., to an induced sub-poset isomorphic to the leftmost poset in Figure 1. Otherwise, the poset $P$ contains an induced sub-poset isomorphic to the second one in Figure 1, and $P$ does not have the fixed point property [15, Ex. 4-7].
@@
Let $P$ be a poset in which every edge of every crown in $E(P)$ belongs to an improper 4-crown. There exists a poset $R$ of height two with the following property: A 4-crown $C\subseteq E(P)$ is a retract of $P$ iff it is a retract-crown of $R$, too.
@@
According to Rival [9], [15, Theorem 4.37], $E(P)$ is dismantlable to a singleton by a sequence of I-retractions, retracting in the first step, say, $\ell\in L(P)$ to $u\in U(P)$. We have ${\uparrow\,}_{E(P)}\ell=\{u\}$ due to the definition of an I-retraction, hence $({\uparrow\,}_{P}\ell)\cap U(P)=\{u\}$, and $P$ is I-dismantlable to $P^{\prime}:=P|_{X\setminus\big{(}[\ell,u]_{P}\setminus\{u\}\big{)}}$ according to [15, Ex. 4-5]. Due to $E(P^{\prime})=E(P)\setminus\{\ell\}$ we can continue in processing the sequence of I-retractions of $E(P)$ in this way, and everything is shown.

### v6

@@ 1 Introduction
With the exception of Corollary 4, all our results are valid also for infinite posets without infinite chains. Even if the maximal sets addressed in Definition 3 do not exist, everything remains correct if we replace ${\cal F}(P)$ by ${\cal F}_{0}(P)$ in Definition 4 of the multigraph $\mathfrak{F}(P)$.
@@ 2.1 Notation and basic terms
Due to $[a,v]_{P}\cap[b,w]_{P}=[a,w]_{P}\cap[b,v]_{P}$, this definition is independent of the choice of the disjoint edges of $C$. As indicated by the addendum “in $P$”, a 4-crown is proper or improper only with respect to the poset it belongs to, not as an induced sub-poset as such. However, because the ambient poset is always clear in what follows, we omit explicit reference to it in most cases. In parts of the early literature [5, 6, 8, 10, 12, 14], improper 4-crowns are not regarded as crowns.
@@
For a homomorphism $f$ between posets or graphs, we denote by $f|_{Z}$ the pre-restriction of $f$ to a subset $Z$ of its domain. An order homomorphism $r:P\rightarrow P$ is called a retraction of the poset $P$ iff $r$ is idempotent, and an induced sub-poset $R$ of $P$ is called a retract of $P$ iff a retraction $r:P\rightarrow P$ exists with $r[X]=R$. For the sake of simplicity, we identify $r$ with its post-restriction and write $r:P\rightarrow R$. A poset $P$ has the fixed point property iff every retract of $P$ has the fixed point property [16, Theorem 4.8]. An improper 4-crown cannot be a retract of $P$.
@@ 2.2 Results about homomorphisms
Let $C\subseteq E(P)$. There exists a retraction $f:P\rightarrow C$ iff there exists a partition $A,B$ of $L(P)$ and a partition $V,W$ of $U(P)$ fulfilling
$\begin{split}&\>\textrm{if }\>F\subseteq E(P)\>\textrm{is a four-element set with}\>F\cap N\not=\emptyset\>\textrm{ for all}\>\\ &\;N\in\{A,B,V,W\},\>\textrm{then}\>F\>\textrm{is not an improper 4-crown of }\>P\end{split}$ (3)
in which the points $a,b,v,w$ all belong to different sets.
@@
$\Rightarrow$” holds because the existence of a retraction $f:P\rightarrow C$ implies the existence of a retraction of $P$ onto $C$ which is strict on $E(P)$.
$\Leftarrow$”: Assume that there exists a partition $A,B$ of $L(P)$ and a partition $V,W$ of $U(P)$ with the specified properties. Let $f:E(P)\rightarrow C$ be a strict surjective homomorphism induced by these partitions.
For $x\in X$, we define the sets $\alpha(x)$ and $\beta(x)$ as in Proposition 1. Then $\alpha(x)\subseteq\{a,b\}$ and $\beta(x)\subseteq\{v,w\}$. Assume $\alpha(x)=\{a,b\}$ and $\beta(x)=\{v,w\}$. Then $x\in M(P)$, and there exist points $a^{\prime}\in f^{-1}(a)\cap{\downarrow\,}_{P}x$, $b^{\prime}\in f^{-1}(b)\cap{\downarrow\,}_{P}x$, $v^{\prime}\in f^{-1}(v)\cap{\uparrow\,}_{P}x$, and $w^{\prime}\in f^{-1}(w)\cap{\uparrow\,}_{P}x$. But then $F:=\{a^{\prime},b^{\prime},v^{\prime},w^{\prime}\}\subseteq E(P)$ is an improper 4-crown of $P$ with $F\cap N\not=\emptyset$ for all $N\in\{A,B,V,W\}$ in contradiction to (3). At least one of the sets $\alpha(x)$ and $\beta(x)$ is thus a singleton, and Proposition 1 yields the existence of an extension $g:P\rightarrow C$ of $f$. With an appropriate automorphism $\pi$ of $C$, the mapping $\pi\circ g$ is a retraction.
@@ 3.1 The multigraphs $\mathfrak{C}$ and $\mathfrak{F}(P)$
Let $x,x^{\prime}\in L(P)$ and $y,y^{\prime}\in U(P)$ with $C^{\prime}=\{x,x^{\prime},y,y^{\prime}\}$ being a 4-crown. A homomorphism $\phi:\mathfrak{F}(P)\rightarrow\mathfrak{C}$ is called $C^{\prime}$-separating iff
for all $\>\textrm{for all }\>F\in{\cal F}(P)\>\textrm{ :}\>$$\begin{cases}\;x\in F&\quad\Rightarrow\quad\phi(F)\notin{\cal A},\\ \;x^{\prime}\in F&\quad\Rightarrow\quad\phi(F)\notin{\cal B},\\ \;y\in F&\quad\Rightarrow\quad\phi(F)\notin{\cal V},\\ \;y^{\prime}\in F&\quad\Rightarrow\quad\phi(F)\notin{\cal W}.\end{cases}$ (4)
@@
A 4-crown $C\subseteq E(P)$ is a retract of $P$ iff a $C$-separating homomorphism from $\mathfrak{F}(P)$ to $\mathfrak{C}$ exists.
@@
$\Rightarrow$”: Assume that $C$ is a retract of $P$. According to Lemma 1, there exists a partition $A,B$ of $L(P)$ and a partition $V,W$ of $U(P)$ fulfilling (3) in which the points of $C$ all belong to different sets. With $g:E(P)\rightarrow C$ being the strict retraction induced by these partitions, it is easily seen that the mapping
@@
is a homomorphism from $\mathfrak{F}(P)$ to $\mathfrak{C}$. We have $g(a)=a$, thus $a\in F\Rightarrow\phi(F)\notin{\cal B}$. With the corresponding results for the other points of $C$ we see that $\phi$ is $C$-separating.
$\Leftarrow$”: Let $\phi:\mathfrak{F}(P)\rightarrow\mathfrak{C}$ be a $C$-separating homomorphism. We define
@@
Because $\phi$ is $C$-separating, we cannot have $a,b\in A_{0}$ or $a,b\in B_{0}$. Regardless of which of the points $a$ and $b$ belong to $L(P)\setminus(A_{0}\cup B_{0})$ (none, one, or both), we can add them (if required) to the sets $A_{0}$ and $B_{0}$ in such a way that each of the resulting sets $A_{1}\supseteq A_{0}$ and $B_{1}\supseteq B_{0}$ contains exactly one of them. Now we distribute the elements of $L(P)\setminus(A_{1}\cup B_{1})$ arbitrarily on the sets $A_{1}$ and $B_{1}$. The resulting sets $A\supseteq A_{1}$ and $B\supseteq B_{1}$ form a partition of $L(P)$ with $a$ and $b$ belonging to different sets. In the same way, we construct a partition $V,W$ of $U(P)$ with $V_{0}\subseteq V$ and $W_{0}\subseteq W$ in which $V$ and $W$ both contain exactly one of the points $v$ and $w$.
Let $G\in{\cal F}(P)$. There exists a set ${\cal N}\in\{{\cal A},{\cal B},{\cal V},{\cal W}\}$ with $\phi(G)\in{\cal N}$, say, ${\cal N}={\cal A}$. Then $L(P)\cap G\subseteq A_{0}$, thus $G\cap B=\emptyset$. In the same way, we see that also for the other choices of ${\cal N}$ there exists a set $N\in\{A,B,V,W\}$ with $G\cap N=\emptyset$, and Lemma 1 delivers the existence of a retraction of $P$ onto $C$.
@@
Let $Z:=\left(\cup\;{\cal F}(P)\right)\cup{\cal M}(P)\cup C$ and $Q:=P|_{Z}$. Because of $\mathfrak{F}(P)=\mathfrak{F}(Q)$ and Theorem 1, the 4-crown $C$ is a retract of $P$ iff it is a retract of $Q$. The elements of $X\setminus Z$ are thus irrelevant for $C$ being a retract of $P$ or not. And indeed, for $x\in M(P)\setminus Z$, one of the sets $L(P)\cap{\downarrow\,}x$ or $U(P)\cap{\uparrow\,}x$ is a singleton, and according to [16, Ex. 4-5], the point $x$ can be removed by a series of I-retractions. By doing so for all points in $M(P)\setminus Z$, we arrive at the poset $P^{\prime}:=P|_{E(P)\cup Z}$. The poset $P^{\prime}|_{Z}$ contains $C$, the improper 4-crowns of $P$ and their inner, and the poset $P^{\prime}|_{E(P)\setminus Z}$ has height one or is an antichain. Intuitively, it is plausible that a homomorphism mapping $P^{\prime}|_{Z}$ onto $C$ can be extended to the total poset $P^{\prime}$.
@@ 4 Additional results about crowns as retracts
For the next theorem, we need a slight variant of Rival [12, pp.  118-119], [16, Lemma 4.35]. The proof is a copy of the short proof of [16, Lemma 4.35]:
@@
We conclude that for $P$ having the fixed point property, every edge of every crown in $E(P)$ must belong to an improper 4-crown in $E(P)$. (Examples are the rightmost poset in Figure 1 and the posets in Figure 3.) For a poset $P$ of height two, this necessary condition can be sharpened: every edge of every crown in $E(P)$ must belong to an hourglass-crown, i.e., to an induced sub-poset isomorphic to the leftmost poset in Figure 1. Otherwise, the poset $P$ contains an induced sub-poset isomorphic to the second one in Figure 1, and $P$ does not have the fixed point property [16, Ex. 4-7].
@@
Let $P$ be a poset in which every edge of every crown in $E(P)$ belongs to an improper 4-crown. There exists a poset $R$ of height two with the following properties:
  • The 4-crown bundles of $R$ are exactly the 4-crown bundles of $P$.
  • A 4-crown $C\subseteq E(P)$ is a retract of $P$ iff it is a retract-crown of $R$, too.
@@
According to Rival [10], [16, Theorem 4.37], $E(P)$ is dismantlable to a singleton by a sequence of I-retractions, retracting in the first step, say, $\ell\in L(P)$ to $u\in U(P)$. We have ${\uparrow\,}_{E(P)}\ell=\{\ell,u\}$ due to the definition of an I-retraction, hence $({\uparrow\,}_{P}\ell)\cap U(P)=\{u\}$, and $P$ is I-dismantlable to $P^{\prime}:=P|_{X\setminus\big{(}[\ell,u]_{P}\setminus\{u\}\big{)}}$ according to [16, Ex. 4-5]. Due to $E(P^{\prime})=E(P)\setminus\{\ell\}$ we can continue in processing the sequence of I-retractions of $E(P)$ in this way, and everything is shown.